# Networks of neural networks: more is different

### v1

@@ 1 Introduction
The celebrated constructive criticism to the reductionist hypothesis more is different – a concept popularized by Philip W. Anderson in the 70’s AndersonScience – is a foundational statement in Statistical Mechanics and its manifestations are ubiquitous in Nature, from phase transitions in Physics Derrida1984 ; Goldenfeld2018 and Chemistry FrenkelSmith ; EyesChem to collective behaviors in Biology CavagnaCabibbo ; Bialek2012 and Ecology Sole2006 ; Chiara2021 (see also Parisi’s Nobel Prize Lecture GiorgioNobelLecture ). In this paper, we inspect this principle at work with Hopfield associative neural networks Hopfield , each of which, independently, can perform only a specific task, that is, pattern recognition Bishop .
@@
In particular, we consider an ensemble of Hopfield networks that share the same dataset of random, binary patterns Coolen and couple them through repulsive interactions. Our findings demonstrate that the resulting network of networks can execute tasks that exceed the capabilities of any single constituting network. Specifically, the combined system exhibits the ability to perform pattern disentanglement —i.e., when presented with a mixture of patterns, it can separate the input into the original components. In fact, a composite system of, say, $L$ Hopfield networks displays the natural architecture to disentangle combinations of $L$ patterns; the mixtures that we will consider here are obtained by applying a majority rule to $L$ patterns drawn from the dataset, and this produces the so-called spurious states, known to emerge as (unwanted) minima in a single Hopfield model Amit .
@@
It is worth noticing that our assembly of $L$ interacting Hopfield networks can also be looked at as an $L$-directional associative memory Kosko ; BCDS-JPA2023 ; AMT-JSP2018 endowed with Hebbian intra-layer interactions where intra-layer interactions are attractive but inter-layer interactions are repulsive (i.e. their sign is reversed, unlike classic directional associative memories). Without such a reversal, pattern disentanglement would be prevented as layers would tend to align to the same pattern, unless the task is simplified to disentangling mixtures of patterns drawn from independent datasets (a simpler task [^1] In this scenario, mixtures states can not be seen as Hopfield spurious states. that can be handled by standard hetero-associative neural networks AABCR-2024 ).
@@
From a theoretical standpoint, this new capability of the model under study allows for further dissecting the world of spurious states and it may shed further light on the complex landscape of the Hopfield model itself, the latter being a cornerstone for pattern recognition and associative memories, as highlighted by the recent Nobel Prize award in Physics NP . On the practical side, the potential applications are vast. Recalling that the most stable spurious states of the Hopfield model are mixtures built of by triplets of patterns Amit , one can consider, for instance, video signals, where colors emerge from the combination of three primary colors (red, yellow, and blue), or audio signals, where chords consist of three primary notes (as, e.g. the C-Major chord is a triad formed from a root C, a major third E and a perfect fifth G). However, rather than focusing on specific applications, our aim here is to describe the network’s emergent computational properties and uncover the fundamental mechanisms underpinning them, in the context of synthetic datasets.
@@
The paper is structured as follows. In Sec. 2, we introduce the model and we present the main analytical results obtained by employing statistical-mechanics tools. In Sec. 3, we explore its ability to perform pattern disentanglement by different approaches: in Sec. 3.1, we study analytically the stability of several paradigmatic configurations, e.g., where each layer is aligned with the $L$-pattern mixtures (playing as the network input) and where each layer is aligned with a different pattern participating in the mixture (playing as the target network output); next, in Sec. 3.2, we make this analysis more accurate by examining the sign of the free-energy Hessian matrix, whence we get insights on the stability of the input and of the target configurations; then, in Sec. 3.3, we proceed the investigation by finding a numerical solution of the self-consistency equations stemming from the statistical-mechanics analysis and by suitably revising the standard protocols designed to check retrieval capabilities in order to account for the disentanglement task; finally, in Sec. 3.4, the previous theoretically-driven results are corroborated by Monte Carlo (MC) simulations. This thorough analysis suggests adjustments to the model that could enhance its performance, which will be discussed in Sec. 4. In the final Sec. 5, we summarize results and discuss some outlooks. Technical details on analytical computations are collected in the Appendices A-B. Moreover, in Appendix C we check the robustness of the results by running analogous experiments, but setting $L=5$.
2 Hebbian networks of networks
@@
In this section we introduce the general model, whose architecture is sketched in Fig. 1; to avoid ambiguities, we will refer to a single Hopfield network as a layer. Thus, let us consider $L$ layers, each composed of $N$ binary neurons, denoted as $\bm{\sigma}^{a}=(\sigma^{a}_{1},...,\sigma^{a}_{N})\in\{-1,+1,\}^{N}$ for $a=1,...,L$, that interact pairwise as specified by the following Hamiltonian (or energy or cost function):
@@
Figure 1: Schematic representation of the model under study for the case $L=3$. The three contributions making up the Hamiltonian (5) are highlighted: imitative intra-layer interactions, anti-imitative inter-layer interactions and the coupling with an external field.
where $\bm{\xi}^{\mu}=(\xi^{\mu}_{1},...,\xi^{\mu}_{N})\in\{-1,+1\}^{N}$ is the $\mu$-th pattern for $\mu=1,...,K$ and $\bm{g}\in\mathbb{R}^{L\times L}$ specifies the nature (imitative or anti-imitative) of intra-layer and inter-layer interactions. By introducing the Mattis magnetization
@@
that assesses the retrieval of the $\mu$-th pattern by the $a$-th layer, we can recast (1) as
@@
$g_{ab}=\begin{cases}\leavevmode\nobreak\ \leavevmode\nobreak\ 1\leavevmode\nobreak\ \textrm{if}\leavevmode\nobreak\ a=b\\ -\lambda\leavevmode\nobreak\ \textrm{if}\leavevmode\nobreak\ a\neq b\end{cases}$ (4)
with $\lambda\in[0,(L-1)^{-1})$ to ensure that $\bm{g}$ is positive definite (vide infra). This implies that neurons belonging to the same layer interact by imitative Hebbian coupling – namely, each layer tends to align to a single pattern, as it is the case in the standard Hopfield model – while neurons belonging to different layers interact by anti-imitative Hebbian coupling – namely, configurations where all layers are aligned with the very same pattern are discouraged, consistently with the kind of task we are interested in. In any case, we stress that the Hebbian shape of the interaction is preserved and, as expected, in the limit $\lambda\to 0$ the model reduces to a collection of $L$ independent Hopfield models trained on the same dataset of patterns $\{\bm{\xi}^{\mu}\}_{\mu=1,..,K}$. The structure of the Hamiltonian (3) resembles that of $L$-directional associative memories Kosko ; Peliti ; BCDS-JPA2023 ; AABCR-2024 , but in those models intra-layer couplings are absent and the inter-layer couplings are imitative.
In general, we can allow for an external field, tuned by the scalar $H\in\mathbb{R}^{+}$ and pointing in the direction specified by $\bm{h}^{a}\in\{-1,+1\}^{N}$ for $a=1,...,L$, namely
@@
Notice the variables and the parameters which the Hamiltonian depends on: beyond the model’s degrees of freedom $\bm{\sigma}$, there appear the external fields $\bm{h}=\{\bm{h}^{a}\}_{a=1,...,L}$ that are quenched and will be chosen according to the application we aim to address with these networks (see Sec. 3 and AABCR-2024 ), the pattern dataset $\bm{\xi}=\{\bm{\xi}^{\mu}\}_{\mu=1,...,K}$, that is quenched and drawn from a Rademacher distribution such that
@@
To see the interplay between the contributions making up the Hamiltonian (5) (we recall that the first two contributions correspond to the sum of $L$ Hopfield models, while the third contribution introduces a coupling among them), let us set $H=0$ and notice that, in order to minimize the first contribution, the neurons in each layer tend to align with an arbitrary pattern, say $\bm{\sigma}^{a}=\bm{\xi}^{\mu}$, and, since patterns are (approximately [^2] This follows from the choice (6), from which $\bm{\xi}^{\mu}\cdot\bm{\xi}^{\nu}\approx\delta_{\mu,\nu}$, with negligible corrections in the limit $N\to\infty$.) orthogonal, it follows that $m^{a}_{\mu}=1$ and $m^{a}_{\nu}=0$ for $\nu\neq\mu$; in order to minimize the third contribution, the pattern retrieved by different layers must be the same, apart from the sign [^3] This intrinsic blemish will be fixed in Sec. 4 by adopting higher-order inter-later interactions, in such a way that the third contribution will as well favor the disentangled state.: assuming $L$ even, $L/2$ layers are aligned with $\bm{\xi}^{\mu}$ and the other $L/2$ layers are aligned with $-\bm{\xi}^{\mu}$ in such a way that $\sum_{a\neq b}m^{a}_{\mu}m^{b}_{\mu}=-L/2$ (when $L$ is odd, the unbalance makes the sum equal to $(L-1)/2$). Notice that the case where $\bm{\sigma}^{a}=\bm{\xi}^{\mu}$ and $\bm{\sigma}^{b}=\bm{\xi}^{\nu}$, with $\nu\neq\mu$ if $b\neq a$, minimizes the first contribution but is only a local minimum for the third contribution, which would approximately [^4] Again, this follows from the choice (6). equal zero.
The statistical-mechanics investigation of the model is detailed in the App. A by exploiting interpolating techniques (see e.g., guerra2001sum ; Agliari-NN2020 ), while here we simply report the explicit expression of the quenched free energy $\mathcal{A}^{RS}$ found in the thermodynamic limit $N\to\infty$, under the replica-symmetry (RS) approximation and in the high-storage regime. Before presenting it, we anticipate that, beyond the Mattis magnetizations $\bm{m}^{a}$, for $a=1,...,L$, another set of macroscopic observables needs to be defined, that is,
@@
which represents the overlap between the neural configurations of two replicas ${\bm{\sigma}}^{a,(1)}$ ${\bm{\sigma}}^{a,(2)}$, where the superscripts $(1)$ and $(2)$ denote the replica index. The above-mentioned RS approximation implies that, in the thermodynamic limit, the distribution of these macroscopic observables concentrates around their expectation values denoted as, respectively, $\bar{m}^{a}_{\mu}$ and $\bar{q}_{12}^{a}$ for $\mu=1,...,K$ and $a=1,...,L$.
@@
where $\gamma=\lim_{N\to\infty}K/N$ defines the storage capacity in the network,
$\begin{array}[]{lll}\bar{p}_{12}^{a}&=&-\displaystyle\sum\limits_{b\neq a}^{L}\sqrt{\dfrac{\bar{q}_{12}^{b}}{\bar{q}_{12}^{a}}}\;(\mathcal{G}^{-1}\,)_{ab}\,-\displaystyle\sum\limits_{c,b=1}^{L}\sqrt{\bar{q}_{12}^{c}\bar{q}_{12}^{b}}\;\left[\partial_{\bar{q}_{12}^{a}}(\mathcal{G}^{-1}\,)_{cb}\,\right],\end{array}$ (9)
@@
Although these expressions look fairly standard, when the expectation $\mathbb{E}_{\bm{\xi},x}$ is implemented, they become rather cumbersome, see for instance App. A and App. B. For this reason, their numerical solution will be limited to the low-load regime ($\gamma=0$), see Sec. 3.3.

### v2

# Networks of Hebbian networks: more is different

@@ 1 Introduction
The celebrated constructive criticism to the reductionist hypothesis more is different – a concept popularized by Philip W. Anderson in the 70’s AndersonScience – is a foundational statement [^1] The phrase emphasizes that, in a complex network, collective phenomena emerge that cannot be predicted from – or reduced to – the behavior of its isolated nodes: for instance, the Hopfield model is able to recognize a pattern due to the inner dialogues among its neurons (i.e. the nodes of the network) but none of them has even the concept of pattern, thus the question addressed in this manuscript: are new collective phenomena appearing by constructing networks of Hopfield models? in Statistical Mechanics and its manifestations are ubiquitous in Nature, from phase transitions in Physics Derrida1984 ; Goldenfeld2018 and Chemistry FrenkelSmith ; EyesChem to collective behaviors in Biology CavagnaCabibbo ; Bialek2012 and Ecology Sole2006 ; Chiara2021 . In this paper, we inspect this principle at work with Hopfield associative neural networks Hopfield , each of which, independently, can perform only a specific task, that is, pattern recognition Bishop .
@@
In particular, we consider an ensemble of Hopfield networks that share the same dataset of random, binary patterns Coolen and couple them through repulsive interactions. Our findings demonstrate that the resulting network of networks can execute tasks that exceed the capabilities of any single constituting network. Specifically, the combined system exhibits the ability to perform pattern disentanglement —i.e., when presented with a mixture of patterns, it can separate the input into the original components. In fact, a composite system of, say, $L$ Hopfield networks displays the natural architecture to disentangle combinations of $L$ patterns; the mixtures that we will consider here are obtained by applying a majority rule to $L$ patterns drawn from the dataset: this produces the so-called spurious states, known to emerge as (unwanted) minima in a single Hopfield model Amit .
@@
It is worth noticing that our assembly of $L$ interacting Hopfield networks can also be looked at as an $L$-directional associative memory Kosko ; BCDS-JPA2023 ; AMT-JSP2018 endowed with Hebbian interactions where intra-layer interactions are attractive but inter-layer interactions are repulsive (i.e. their sign is reversed, unlike classic directional associative memories). Without such a reversal, pattern disentanglement would be prevented as layers would tend to align to the same pattern, unless the task is simplified to disentangling mixtures of patterns drawn from independent datasets (a simpler task [^2] In that scenario, mixtures states can not be seen as Hopfield spurious states. that can be handled by standard hetero-associative neural networks AABCR-2024 ).
@@
From a theoretical standpoint, this new capability of the model under study allows for further dissecting the world of spurious states and it may shed further light on the complex landscape of the Hopfield model itself. On the practical side, the potential applications are vast. Recalling that the most stable spurious states of the Hopfield model are mixtures built of by triplets of patterns Amit , one can consider, for instance, video signals, where colors emerge from the combination of three primary colors (red, yellow, and blue), or audio signals, where chords consist of three primary notes (as, e.g. the C-Major chord is a triad formed from a root C, a major third E and a perfect fifth G). However, rather than focusing on specific applications, our aim here is to construct a quantitative theory able to describe the network’s emergent computational properties and uncover the fundamental mechanisms underpinning them, in the context of synthetic datasets.
It is worth pointing out that, at present, there are already several algorithmic approaches to pattern disentanglement, yet nor any of these is based on Hebbian learning, neither any of these provides a theoretical explanation for this type of information processing by neural networks. For instance, Disentangled Representation Learning DRL2024 plays a central role when the network fails to infer the correct features (as often they have to be disentangled at first) and this, in turn, turned pivotal for explainable AI in visual recognition (e.g., when an image contains several objects to be recognized). Typically, the underlying neural architectures are deep learning scaffolds (namely long multi-layered networks or deep Boltzmann machines) Cacao20 where, as the disentanglement goes by, inner layers naturally split into sub-architectures specialized in the recognition of a specific feature or pattern: Deep Hierarchical Representations DipLer16 fall in this ensemble too.
@@
The paper is structured as follows. In Sec. 2 we introduce the model and Sec. 3 we present the main analytical results obtained by employing statistical-mechanics tools. Next, focusing on the test-case $L=3$, in Sec. 4, we explore its ability to perform pattern disentanglement by different approaches. In Sec. 4.1, we study analytically the stability of several paradigmatic configurations, e.g., where each layer is aligned with the $L$-pattern mixtures and where each layer is aligned with a different pattern participating in the mixture: these two configurations play, respectively, as the network input and the target network output. Next, in Sec. 4.2, we make this analysis more accurate by examining the sign of the free-energy Hessian matrix, whence we get insights on the stability of the input and of the target configurations; then, in Sec. 4.3, we proceed the investigation by finding numerical solutions of the self-consistency equations stemming from the statistical-mechanics analysis and by suitably revising the standard protocols designed to check retrieval capabilities in order to account for the disentanglement task; finally, in Sec. 4.4, the previous theoretically-driven results are corroborated by Monte Carlo (MC) simulations. In the final Sec. 5, we summarize results and discuss some outlooks. Technical details on analytical computations are collected in the Appendices A-D. Moreover, in Appendix E we describe the methodology underlying computational experiments and in Appendix F we check the robustness of the results by running analogous experiments, but setting $L=5$. Next, possible adjustments to the model that could enhance its performance are discussed in Appendix G. Finally, in Appendix H, we show how our theory for pattern disentanglement by the present Hebbian network can also shed light on pattern disentanglement by already existing architectures.
2 Definitions
@@
In this section we introduce the general model, whose architecture is sketched in Figure 1; to avoid ambiguities, we will refer to a single Hopfield network as a layer. Thus, let us consider $L$ layers, each composed of $N$ binary neurons, denoted as $\bm{\sigma}^{a}=(\sigma^{a}_{1},...,\sigma^{a}_{N})\in\{-1,+1,\}^{N}$ for a=1,,La=1,...,L, that interact pairwise as specified by the following cost function (or energy or Hamiltonian):
@@
Figure 1: Schematic representation of the model under study, where we set $L=3$. The three contributions making up the cost function (5) are highlighted: imitative intra-layer interactions (represented by a $\oplus$ loop), anti-imitative inter-layer interactions (represented by a $\lambda$ double arrow) and the coupling with an external field (represented by a single arrow $\bm{h}$).
where $\bm{\xi}^{\mu}=(\xi^{\mu}_{1},...,\xi^{\mu}_{N})\in\{-1,+1\}^{N}$ is the $\mu$-th pattern for $\mu=1,...,K$ and $\bm{g}\in\mathbb{R}^{L\times L}$ specifies the nature (imitative or anti-imitative) of the Hebbian intra-layer and inter-layer interactions. By introducing the Mattis magnetizations
@@
as order parameters that assesses the retrieval of the $\mu$-th pattern by the $a$-th layer [^3] Note that ech $m_{\mu}^{a}\in[-1,+1]$ such that a value of $m_{\mu}^{a}$ close to +1+1 accounts for a retrieved pattern (likewise $m_{\mu}^{a}$ close to 1-1 accounts the retrieval of the inverse pattern $\xi^{\mu}$) while $m_{\mu}^{a}\sim 0$ implies no retrieval., we can recast (1) as
@@
$g_{ab}=\begin{cases}\penalty 10000\ \penalty 10000\ 1\penalty 10000\ \textrm{if}\penalty 10000\ a=b\\ -\lambda\penalty 10000\ \textrm{if}\penalty 10000\ a\neq b\end{cases}$ (4)
with $\lambda\in[0,(L-1)^{-1})$ to ensure that $\bm{g}$ is positive definite (vide infra). This implies that neurons belonging to the same layer interact by imitative Hebbian coupling – namely, each layer tends to align to a single pattern, as it is the case in the standard Hopfield model – while neurons belonging to different layers interact by anti-imitative Hebbian coupling – namely, configurations where all layers are aligned with the very same pattern are discouraged, consistently with the kind of task we are interested in. In any case, we stress that the Hebbian shape of the interaction is preserved and, as expected, in the limit $\lambda\to 0$ the model reduces to a collection of $L$ independent Hopfield models trained on the same dataset of patterns $\{\bm{\xi}^{\mu}\}_{\mu=1,..,K}$. The structure of the cost function (3) resembles that of $L$-directional associative memories Kosko ; Peliti ; BCDS-JPA2023 ; AABCR-2024 , but in those models intra-layer couplings are absent and the inter-layer couplings are imitative. A modular organization of recurrent associative memories (but devoid of intra-layer interactions and exhibiting a hierarchical structure) has been proposed also in DmitryK2021 , inspired by cortical feedback structures, however that kind of network is most suitable for retrieving multiple objects within the same image and not for disentangling spurious mixtures as those faced here.
In general, we can allow for an external field, tuned by the scalar $H\in\mathbb{R}^{+}$ and pointing in the direction specified by $\bm{h}^{a}\in\{-1,+1\}^{N}$ for a=1,,La=1,...,L, namely
@@
Notice the variables and the parameters which the cost function depends on: beyond the model’s degrees of freedom $\bm{\sigma}$, there appear the external fields $\bm{h}=\{\bm{h}^{a}\}_{a=1,...,L}$ that are quenched and will be chosen according to the application we aim to address with these networks [^4] We anticipate that, in this framework, we aim to assess the capacity of the model to tease apart the patterns appearing in mixtures like $\textrm{sign}(\sum_{\mu=1}^{L}\bm{\xi}^{\mu})$, as these are supplied as input. Thus, a natural choice for the external field is precisely $\bm{h}^{a}=\textrm{sign}(\sum_{\mu=1}^{L}\bm{\xi}^{\mu})$, for a=1,,Ma=1,...,M, as this is the information at hand when setting the machine, see also Sec. 4 and AABCR-2024 ; AFM2025 ., the pattern dataset $\bm{\xi}=\{\bm{\xi}^{\mu}\}_{\mu=1,...,K}$, that is quenched and drawn from a Rademacher distribution such that
@@
To see the interplay between the contributions making up the cost function (5) (we recall that the first two contributions correspond to the sum of $L$ Hopfield models, while the third contribution introduces a coupling among them), let us set $H=0$ and notice that, in order to minimize the first contribution, the neurons in each layer tend to align with an arbitrary pattern, say $\bm{\sigma}^{a}=\bm{\xi}^{\mu}$, and, since patterns are (approximately [^5] This follows from the choice (6), from which ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}N^{-1}}\bm{\xi}^{\mu}\cdot\bm{\xi}^{\nu}\approx\delta_{\mu,\nu}$, with negligible corrections in the limit $N\to\infty$.) orthogonal, it follows that $m^{a}_{\mu}=1$ and $m^{a}_{\nu}=0$ for $\nu\neq\mu$; in order to minimize the third contribution, the pattern retrieved by different layers must be the same, apart from the sign [^6] This intrinsic blemish will be fixed in Sec. G by adopting higher-order inter-later interactions, in such a way that the third contribution will as well favor the disentangled state.: assuming $L$ even, $L/2$ layers are aligned with $\bm{\xi}^{\mu}$ and the other $L/2$ layers are aligned with $\bm{\xi}^{\mu}$ in such a way that $\sum_{a=1}^{L}\sum_{b=1,b\neq a}^{L}m^{a}_{\mu}m^{b}_{\mu}=-L/2$ (when $L$ is odd, the unbalance makes the sum equal to (L1)/2(L-1)/2). Notice that the case where $\bm{\sigma}^{a}=\bm{\xi}^{\mu}$ and $\bm{\sigma}^{b}=\bm{\xi}^{\nu}$, with $\nu\neq\mu$ if $b\neq a$, minimizes the first contribution but is only a local minimum for the third contribution, which would approximately [^7] Again, this follows from the choice (6). equal zero.
3 Stationary state description
Before specifying the task that we aim to address with the model introduced above, we carry out a statistical mechanics investigation of the network’s computational capabilities in order to get a description of its expected macroscopic behavior, once a stationary state is reached (at a given temperature $T=\beta^{-1}$). This analysis is detailed in the App. A by exploiting interpolating techniques (see e.g., guerra2001sum ; Agliari-NN2020 ), while here we simply report the explicit expression of the quenched free energy $\mathcal{A}^{RS}$ found in the thermodynamic limit $N\to\infty$, under the replica-symmetry (RS) approximation and in the high-storage regime. Before presenting it, we anticipate that, beyond the Mattis magnetizations $\bm{m}^{a}$, for a=1,,La=1,...,L, another set of macroscopic observables needs to be defined, that is,
@@ 3 Stationary state description
which represents the overlap between the neural configurations of two replicas ${\bm{\sigma}}^{a,(1)}$ ${\bm{\sigma}}^{a,(2)}$, where the superscripts $(1)$ and $(2)$ denote the replica index. The above-mentioned RS approximation implies that, in the thermodynamic limit, the distribution of these macroscopic observables concentrates around their expectation values denoted as, respectively, $\bar{m}^{a}_{\mu}$ and $\bar{q}_{12}^{a}$ for $\mu=1,...,K$ and a=1,,La=1,...,L.
@@
where $\gamma=\lim_{N\to\infty}K/N$ defines the network storage,
$\bar{p}_{12}^{a}=-\displaystyle\sum\limits_{\begin{subarray}{c}b=1\\ b\neq a\end{subarray}}^{L}\sqrt{\dfrac{\bar{q}_{12}^{b}}{\bar{q}_{12}^{a}}}\;(\mathcal{G}^{-1}\,)_{ab}\,-\displaystyle\sum\limits_{c,b=1}^{L}\sqrt{\bar{q}_{12}^{c}\bar{q}_{12}^{b}}\;\left[\partial_{\bar{q}_{12}^{a}}(\mathcal{G}^{-1}\,)_{cb}\,\right],$ (9)
@@
Although these expressions look fairly standard, when the expectation $\mathbb{E}_{\bm{\xi},x}$ is implemented, they become rather cumbersome, see for instance App. A and App. B. For this reason, their numerical solution will be limited to the low-storage regime ($\gamma=0$), see Sec. 4.3.